# Learning-Based Difficulty Calibration for Enhanced Membership Inference Attacks

### v1

@@ 3.3 Attack Framework
Based on the aforementioned intuitions, we propose to build a classifier that can determine the membership of data records. Our ultimate goal is to utilize the discriminatory abilities of neural network models to establish the optimal threshold for various classes based on the membership scores and neighborhood information. The proposed attack consists of two phases. The first phase is the training phase, in which a shadow target model $h$, a reference model $g$, and an attack classifier are all trained. The second phase is the inference phase, in which we obtain the features of each target data record from the trained reference model and the target model $f$ and use the features for classification using the attack classifier. The proposed attack workflow is illustrated in Figure 6.
In the proposed attack, we use an auxiliary dataset $\mathcal{D}_{aux}$ that has the same distribution as the data used for training the target model. During the training phase, we first split $\mathcal{D}_{aux}$ into three distinct parts. 1. $\mathcal{D}^{train}_{shadow}$, which is used to train the shadow target model. 2. $\mathcal{D}^{heldout}_{shadow}$, which contains non-member data records for the shadow model. 3. $\mathcal{D}^{train}_{ref}$, which is used to train the reference model. This way, we can keep a clear separation between the data used for training the different models. The shadow target model should behave similarly to the target model. To achieve this, we train the shadow target model until its validation loss is close to that of the target model. Once all the models are trained, we feed all the data records in $\mathcal{D}^{train}_{shadow}$ and $\mathcal{D}^{heldout}_{shadow}$ to the shadow target model and the reference model.
We can categorize the membership scores obtained from the shadow target model into four groups: members’ $s(h,(x,y))$, non-members’ $s(h,(x,y))$, members’ $s(g,(x,y))$, and non-members’ $s(g,(x,y))$. Note that the data records in $\mathcal{D}^{train}_{shadow}$ and $\mathcal{D}^{heldout}_{shadow}$ are made up of members and non-members of the training dataset of the shadow target model, respectively. This means that the attacking classifier can observe how both members and non-members behave on the shadow target model and the reference model, allowing it to discriminate different behaviors. The reference model is introduced to calibrate the membership scores. Therefore, the membership scores of members on the shadow target model are paired with those on the reference model, and Eq (8) is used to calculate the calibrated membership score $S^{cal}(h,g,(x,y))$. To obtain the neighborhood information of a data record, we exclude the interference of the training data of the reference model by using the data records from $D^{train}_{shadow}$ and $D^{heldout}_{shadow}$ only. These data records consist of $v_{aux}$ in Eq (7). Then, the same process is carried out for non-members to obtain the calibrated membership score.
@@
During the inference phase, an attacker can only access the target model as a black box. This means that they can only access the label and membership score of a target data record $d_{target}$ by using the prediction results of the target model. Just like in the training phase, we use $D^{train}_{shadow}$ and $D^{heldout}_{shadow}$ to obtain neighborhood information. Then, we provide $d_{target}$ to the same reference model that was used during the training phase and calculate the calibrated membership score $S^{cal}(f,g,(x,y))$. Finally, we feed all three values to the classifier to predict the membership of $d_{target}$. Our classifier is an MLP model that consists of two hidden layers with ReLU activation functions, followed by a softmax layer.
@@ 4.1.1 Datasets
In our evaluations, we divide each dataset into five parts: $D_{target}^{train}$, $D_{target}^{heldout}$, $D_{shadow}^{train}$, $D_{shadow}^{heldout}$, and $D_{ref}^{train}$. $D_{target}^{train}$ and $D_{target}^{heldout}$ are utilized as the training and testing dataset for the target model. $D_{shadow}^{train}$ and $D_{shadow}^{heldout}$, on the other hand, are used to train and test the shadow target model. $D_{ref}^{train}$ is the training dataset for the reference model. In our experiments, the datasets are proportionally sized in the ratio of 5:5:3:3:4.
@@ 4.1.2 Models
When selecting target models, we prioritize those with better classification accuracy. For the CIFAR-10 and CINIC-10 datasets, we choose VGG-16 as the target model, while for the CIFAR-100 dataset, we use SmallNet, this is to make a fair comparison with the method by Watson et al. [37] where they use SmallNet too. For the remaining datasets, we employ a multi-layer perceptron (MLP) model as the target model, which consists of one hidden layer with ReLU activation function, followed by a softmax layer. The training and testing accuracy of the target models is shown in Table I.
TABLE I: Accuracy of target models
@@
We use the same network architecture as the target model for both the reference model and the shadow target model. We also explore employing different model architectures for the reference model and shadow target model (results are detailed in Section 5). When training the shadow target model, its validation loss does not need to be similar to the target model (this non-requirement is practically useful since our method doesn’t assume the attacker needs to know the target model training validation loss). The reference model is trained until it reaches the maximum validation accuracy. Our proposed MIA classifier is of multi-layer-perceptron architecture, that consists of two hidden layers with ReLU activation functions, followed by a softmax layer. We utilize stochastic gradient descent (SGD) with a learning rate of 0.1, Nesterov momentum of 0.9, and a cosine learning rate schedule for the training. The duration of training for each model varies between 20 to 200 epochs, depending on the complexity of the models and the size of the datasets.
@@ 4.1.3 Metrics
  • Precision-Recall (PR) Curve In real-world scenarios, attackers are more likely to encounter members that are easy to predict into the model than those that are hard to predict, as indicated in Figure 2. Therefore, most MIAs can achieve high recall rates by identifying such members. However, recall only measures the effectiveness of the model in capturing most of the positive instances, it does not reflect how accurately the model predicts positives. Therefore, it is also important to measure precision values at relatively high recall rates. We can do this by looking at the precision values when the recall value ranges between 0.2 and 0.7. This metric helps us measure how effectively the model balances between precision and recall.
@@ 4.1.4 Baselines
In our evaluations, we compared LDC-MIA with three other MIA methods. Salem et al. [30] used the posteriors of target data records obtained from the target model and trained a shadow model to mimic the target model’s behaviors. They proposed three different adversary models, and we compared LDC-MIA with Adversary 1, which had the best performance. Yeom et al. [39] performed the attack without any auxiliary model by using the loss values of the target data records on the target model. Watson et al. [37] used a reference model for difficulty calibration when performing MIA. To ensure a fair comparison, we used the same auxiliary dataset and two auxiliary models — one shadow model and one reference model — in our proposed attack. For other MIA methods, we use two reference or shadow models if they employ any.
@@ 4.2 Main Results
In this section, we present the results of our proposed attack method and compare it with other MIA methods. All the attacks are carried out in the black-box scenario, and we demonstrate and analyze the results using various metrics as mentioned in Section 4.1. Furthermore, we compare the attack performance and cost of LDC-MIA with LiRA [3], the current state-of-the-art MIA method.
(a) (a) CIFAR-10
(b) (b) CIFAR-100
(c) (b) CINIC-10
(d) (c) Adult
(a) (a) CIFAR-10
(b) (b) CIFAR-100
(c) (b) CINIC-10
(d) (c) Adult
(e) (d) Credit
Figure 7: Full log-scale ROC curves of MIAs on different datasets.
TABLE II: TPR at Low FPR regions of MIAs across datasets across MIA methods.
@@
0.01% 0.1% 1% 0.01% 0.1% 1% 0.01% 0.1% 1% 0.01% 0.1% 1% 0.01% 0.1% 1%
Salem et al. [30] 0.00% 0.1% 1.6% 0.04% 0.3% 2.3% 0.00% 0.2% 2.3% 0.00% 0.1% 1.0% 0.00% 0.0% 0.0%
Yeom et al. [39] 0.01% 0.0% 0.0% 0.04% 0.2% 1.0% 0.00% 0.0% 0.0% 0.00% 0.0% 0.0% 0.50% 0.5% 3.0%
Watson et al. [37] 0.05% 0.5% 3.5% 0.32% 1.1% 5.9% 0.07% 0.4% 4.0% 0.01% 0.2% 2.0% 1.50% 1.5% 4.0%
Ours 0.27% 1.8% 6.4% 0.95% 3.8% 16.4% 0.16% 1.6% 8.7% 0.15% 0.8% 3.5% 2.50% 2.5% 6.5%
4.2.1 TPR at low FPR regions
In this experiment, we compare the TPR-FPR tradeoff of our method LDC-MIA and three other MIA methods across five datasets. The ROC curves for all methods over five datasets are depicted in Figure 7. The figure shows that LDC-MIA achieves higher TPR at almost all FPRs than other MIA methods across all datasets. Further, we compare the TPR values in the low FPR region (i.e., between $0.01\%$ and $1\%$) with results obtained from an average of 5 runs in Table II, The results show that LDC-MIA achieves better TPRs in the low FPR region. On CIFAR-100 and CINIC-10 datasets, LDC-MIA outperforms the state-of-the-art difficulty calibration-based MIA method proposed by Watson et al., having 4x higher TPRs in low FPRs.
(a) (a) CIFAR-10
(b) (b) CIFAR-100
(c) (b) CINIC-10
(d) (c) Adult
(a) (a) CIFAR-10
(b) (b) CIFAR-100
(c) (b) CINIC-10
(d) (c) Adult
(e) (d) Credit
Figure 8: Precision-Recall curves of MIAs on different datasets.
TABLE III: Balanced accuracy and AUC of MIAs on different datasets.
Attack method CIFAR-10 CIFAR-100 CINIC-10 Adult Credit
Balanced acc AUC Balanced acc AUC Balanced acc AUC Balanced acc AUC Balanced acc AUC
Salem et al. [30] 66.30% 0.703 66.54% 0.697 70.14% 0.745 53.59% 0.539 60.50% 0.563
Yeom et al. [39] 69.65% 0.694 84.82% 0.889 77.12% 0.780 55.33% 0.546 60.25% 0.581
Watson et al. [37] 66.88% 0.711 73.78% 0.772 71.80% 0.768 54.90% 0.566 59.25% 0.607
Ours 67.93% 0.754 85.12% 0.918 75.71% 0.832 55.36% 0.577 58.75% 0.608
@@ 4.2.2 Precision-Recall curve
We compare MIA methods’ effectiveness by examining the precision and recall tradeoffs made by our method and three others. The Precision-recall curves of all MIA methods across five datasets are depicted in Figure 8. Our method LDC-MIA reaches the highest precision values over most of the recall value range compared to other MIA methods, across all datasets. For the CIFAR-100 dataset, LDC-MIA achieves a recall rate of $49.1\%$ for members with a precision of $90\%$. For the CIFAR-10 dataset, the recall is $41.7\%$ with a precision of $75\%$. For the CINIC-10 dataset, LDC-MIA identifies $52.72\%$ of the members with a precision of $80\%$.
(a) (a) CIFAR-10
(b) (b) CIFAR-100
(c) (b) CINIC-10
(d) (c) Adult
(a) (a) CIFAR-10
(b) (b) CIFAR-100
(c) (b) CINIC-10
(d) (c) Adult
(e) (d) Credit
@@ 4.2.3 Balanced accuracy and AUC
The balanced accuracy is the arithmetic mean of sensitivity and specificity, the higher the better. AUC value indicates the overall discriminatory power of the model over all possible TPR-FPR tradeoffs, the higher the better. Table III shows the balanced Accuracy and AUC of all the MIA methods averaged out of 5 runs. The highest metric values and the metric values of LDC-MIA have been highlighted. LDC-MIA achieves the highest AUC values across all the datasets. Regarding balanced accuracy values, LDC-MIA has close if not better results compared to the best-performing MIAs.

### v3

@@ 3.3 Attack Framework
Based on the aforementioned intuitions, we propose to build a classifier that can determine the membership of data records. Our ultimate goal is to utilize the discriminatory abilities of neural network models to perform membership inference attack based on the intuitions we identified in Section 3.2. The proposed attack consists of two phases. The first phase is the training phase, in which a shadow target model $h$, a reference model $g$, and an attack classifier are all trained. The second phase is the inference phase, in which we obtain the features of each target data record from the trained reference model and the target model $f$ and use the features for classification using the attack classifier. The proposed attack workflow is illustrated in Figure 6.
In the proposed attack, we use an auxiliary dataset $\mathcal{D}_{aux}$ that has the same distribution as the data used for training the target model. During the training phase, we first split $\mathcal{D}_{aux}$ into three distinct parts. 1. $\mathcal{D}^{train}_{shadow}$, which is used to train the shadow target model. 2. $\mathcal{D}^{heldout}_{shadow}$, which contains non-member data records for the shadow model. 3. $\mathcal{D}^{train}_{ref}$, which is used to train the reference model. This way, we can keep a clear separation between the data used for training the different models.
Training phase. Once all the models are trained, we feed all the data records in $\mathcal{D}^{train}_{shadow}$ and $\mathcal{D}^{heldout}_{shadow}$ to the shadow target model and the reference model. We can then obtain membership scores from both the shadow target model and the reference model on members and non-members, i.e., members’ $s(h,(x,y))$, non-members’ $s(h,(x,y))$, members’ $s(g,(x,y))$, and non-members’ $s(g,(x,y))$. Note that $\mathcal{D}^{train}_{shadow}$ contains data records of the members of the training dataset of the shadow target model, while $\mathcal{D}^{heldout}_{shadow}$ contains data records of non-members. This means that the attacking classifier can observe how both members and non-members behave on the shadow target model and the reference model, allowing it to discriminate different behaviors. The reference model is introduced to calibrate the membership scores. Members’ membership scores on the shadow target model are paired with those on the reference model, and Eq (8) is used to calculate the calibrated membership score $S^{cal}(h,g,(x,y))$. To obtain the neighborhood information of a data record, we exclude the interference of the training data of the reference model by using the data records from $D^{train}_{shadow}$ and $D^{heldout}_{shadow}$ only. These data records consist of $v_{aux}$ in Eq (7), which are obtained from the reference model $g$. Then, the same process is carried out for non-members to obtain the calibrated membership score.
@@
Inference phase. During the inference phase, an attacker can only access the target model as a black box. This means that they can only access the membership score of a target data record $(x_{target},y_{target})$ by using the prediction results of the target model. Just like in the training phase, we use $D^{train}_{shadow}$ and $D^{heldout}_{shadow}$ to obtain neighborhood information. Then, we provide $(x_{target},y_{target})$ to the same reference model that was used during the training phase and calculate the calibrated membership score $S^{cal}(f,g,(x,y))$. Finally, we feed the membership score, calibrated membrship score, ground truth label $y_{target}$ to the classifier to predict the membership of $(x_{target},y_{target})$. Our classifier is an MLP model that consists of two hidden layers with ReLU activation functions, followed by a softmax layer.
@@ 4.1.1 Datasets
In our evaluations, we split each dataset into six parts: $D_{target}^{train}$, $D_{target}^{heldout}$, $D_{shadow}^{train}$, $D_{shadow}^{heldout}$, $D_{ref}^{train}$ and $D^{test}$. $D_{target}^{train}$ is used to train the target model, while $D_{target}^{heldout}$ is made up of non-members of the target model. Similarly, $D_{shadow}^{train}$ is used to train the shadow target model, and $D_{shadow}^{heldout}$ contains non-members for the shadow target model. $D_{ref}^{train}$ is the training dataset for the reference model, and $D^{test}$ is the test dataset for all the models. The sizes of all the datasets used in our experiments are listed in Table I.
TABLE I: Datasets division.
Datasets $D_{target}^{train}$ $D_{target}^{heldout}$ $D_{shadow}^{train}$ $D_{shadow}^{heldout}$ $D_{ref}^{train}$ $D^{test}$
CIFAR-10 12500 12500 7500 7500 10000 10000
CIFAR-100 12500 12500 7500 7500 10000 10000
CINIC-10 22500 22500 13500 13500 18000 90000
Adult 8140 8140 4884 4884 6513 16281
Credit 200 200 120 120 160 200
@@ 4.1.2 Models
To demonstrate the effectiveness of LDC-MIA, we select two models of different sizes as target models on the CIFAR-10 and CIFAR-100 datasets. For the CIFAR-10 dataset, we choose WideResNet28-10[40] and VGG-16[32]. For the CIFAR-100 dataset, we select DenseNet-121[14] and SmallNet. Using SmallNet allows us to make a fair comparison with the existing MIAs. For the CINIC-10 dataset, we choose VGG-16[32]. For the Adult and Credit datasets, we employ a multi-layer perceptron (MLP) model as the target model. This model consists of one hidden layer with ReLU activation function, followed by a softmax layer. The training and testing accuracy of the target models are shown in Table II.
TABLE II: Accuracy of target models
@@
(a) (a) CIFAR-10 (WRN28-10)
(b) (b) CIFAR-10 (VGG-16)
(c) (c) CIFAR-100 (DenseNet-121)
(d) (d) CIFAR-100 (SmallNet)
(e) (e) CINIC-10
(f) (f) Adult
(g) (g) Credit
Figure 7: Full log-scale ROC curves of MIAs(LDC-MIA, Yeom et al. [39], Watson et al. [37] and Salem et al. [29]) on different datasets.
We use the same network architecture as the target model for both the reference model and the shadow target model. We also explore employing different model architectures for the reference model and shadow target model (results are detailed in Section 5). When training the shadow target model, its validation loss does not need to be similar to the target model (this non-requirement is useful in practice since our method does not assume that the attacker know the target model’s validation loss). The reference model is trained until it reaches the maximum validation accuracy. Our proposed MIA classifier is of multi-layer-perceptron architecture, that consists of two hidden layers with ReLU activation functions, followed by a softmax layer. We utilize stochastic gradient descent (SGD) with a learning rate of 0.1, Nesterov momentum of 0.9, and a cosine learning rate schedule for the training. The duration of training for each model varies between 20 to 200 epochs, depending on the complexity of the models and the size of the datasets. All experiments are conducted on general-purpose machines equipped with Intel Xeon Silver 4208 CPU@2.10 GHz, Quadro RTX 5000 GPU, and 16 GB RAM.
@@ 4.1.3 Metrics
  • Precision-Recall (PR) Curve In real-world scenarios, attackers are more likely to encounter members that are easy to predict into the model than those that are hard to predict, as indicated in Figure 2. Therefore, most MIAs can achieve high recall by identifying such members. However, recall only measures the effectiveness of the model in capturing most of the positive instances, it does not reflect how accurately the model predicts positives. Therefore, it is also important to measure precision values at relatively high recall. We can do this by looking at the precision values when the recall value ranges between 0.2 and 0.7. This metric helps us measure how effectively the model balances between precision and recall.
@@ 4.1.4 Baselines
In our evaluations, we compared LDC-MIA with four other MIA methods. Salem et al. [29] used the posteriors of target data records obtained from the target model and trained a shadow model to mimic the target model’s behaviors. They proposed three different adversary models, and we compared LDC-MIA with Adversary 1, which had the best performance. Yeom et al. [39] performed the attack without any auxiliary model by using the loss values of the target data records on the target model. Watson et al. [37] used a reference model for difficulty calibration when performing MIA. LiRA [3] aims to exploit statistical differences between data points labeled as members and non-members to infer membership status. It requires to train a large number of shadow models for each target sample. We compare our work with Salem et al. [29], Yeom et al. [39] and Watson et al. [37] in Section 4.2 and with LiRA [3] in Section 4.3.
@@ 4.2 Main Results
In our evaluations, all the attacks are carried out in the black-box scenario, and we demonstrate and analyze the results using various metrics as mentioned in Section 4.1. To ensure a fair comparison, we used the same auxiliary dataset and two auxiliary models – one shadow model and one reference model – in our proposed attack. For other MIA methods, we use two reference or shadow models if they employ any.
(a) (a) CIFAR-10 (WRN28-10)
(b) (b) CIFAR-10 (VGG-16)
(c) (c) CIFAR-100 (DenseNet-121)
(d) (d) CIFAR-100 (SmallNet)
(e) (e) CINIC-10
(f) (f) Adult
(g) (g) Credit
Figure 8: Precision-Recall curves of MIAs (LDC-MIA, Yeom et al. [39], Watson et al. [37] and Salem et al. [29]) on different datasets.
TABLE III: TPR at Low FPR regions of MIAs across datasets across MIA methods.
Attack Method (FPR) CIFAR-10 CIFAR-100 CINIC-10 Adult Credit
WRN28-10 VGG-16 DenseNet-121 SmallNet
0.01% 0.1% 1% 0.01% 0.1% 1% 0.01% 0.1% 1% 0.01% 0.1% 1% 0.01% 0.1% 1% 0.01% 0.1% 1% 0.01% 0.1% 1%
Salem et al. [29] 0.01% 0.2% 2.5% 0.02% 0.1% 1.2% 0.02% 0.2% 2.3% 0.04% 0.3% 2.3% 0.00% 0.2% 2.3% 0.00% 0.1% 1.0% 0.00% 0.0% 1.5%
Yeom et al. [39] 0.00% 0.0% 0.0% 0.00% 0.0% 0.0% 0.00% 0.0% 2.9% 0.04% 0.2% 1.0% 0.00% 0.0% 0.0% 0.00% 0.0% 0.0% 0.50% 0.5% 3.0%
Watson et al. [37] 0.24% 1.0% 3.3% 0.04% 0.4% 3.4% 0.01% 0.9% 6.5% 0.32% 1.1% 5.9% 0.07% 0.4% 4.0% 0.04% 0.4% 2.0% 1.50% 1.5% 4.0%
Ours 0.66% 2.3% 9.1% 0.21% 1.6% 6.7% 1.28% 4.2% 23.2% 0.95% 3.8% 16.4% 0.16% 1.6% 8.7% 0.20% 1.2% 3.9% 2.50% 2.5% 6.5%
4.2.1 TPR at low FPR regions
In this experiment, we compare the TPR-FPR tradeoff of our method LDC-MIA and three other MIA methods across five datasets. The ROC curves for all methods over five datasets are depicted in Figure 7. The figure shows that LDC-MIA achieves higher TPR at almost all FPRs than other MIA methods across all datasets. Further, we compare the TPR values in the low FPR region (i.e., between $0.01\%$ and $1\%$) with results obtained from an average of 5 runs in Table IV, The results show that LDC-MIA achieves better TPRs in the low FPR region. On CIFAR-100 and CINIC-10 datasets, LDC-MIA outperforms the state-of-the-art difficulty calibration-based MIA method proposed by Watson et al., having 4x higher TPRs in low FPRs.
TABLE IV: Balanced accuracy and AUC of MIAs on different datasets.
@@ 4.2.1 TPR at low FPR regions
WRN28-10 VGG-16 DenseNet-121 SmallNet
Accuracy AUC Accuracy AUC Accuracy AUC Accuracy AUC Accuracy AUC Accuracy AUC Accuracy AUC
Salem et al. [29] 69.41% 0.731 65.19% 0.679 82.28% 0.885 66.54% 0.679 70.14% 0.745 51.17% 0.514 54.75% 0.539
Yeom et al. [39] 68.80% 0.723 69.81% 0.696 82.91% 0.900 84.82% 0.889 77.12% 0.780 55.07% 0.542 60.25% 0.581
Watson et al. [37] 66.20% 0.707 66.18% 0.705 71.10% 0.741 73.78% 0.772 71.80% 0.768 54.86% 0.574 59.25% 0.607
Ours 71.62% 0.794 67.82% 0.752 82.94% 0.933 85.12% 0.918 75.71% 0.832 56.27% 0.592 59.50% 0.640
@@ 4.2.2 Precision-Recall curve
We compare MIA methods’ effectiveness by examining the precision and recall tradeoffs made by our method and three others. The Precision-recall curves of all MIA methods across five datasets are depicted in Figure 8. Note that when the recall is 0, the precision values of all MIAs are also 0 in the figure. Our method LDC-MIA reaches the highest precision values over most of the recall value range compared to other MIA methods, across all datasets. For example, LDC-MIA achieves recall of $77.2\%$ and $49.1\%$ on DenseNet-121 and SmallNet, with $90\%$ precision, respectively for the CIFAR-100 dataset. For the CINIC-10 dataset, LDC-MIA identifies $52.72\%$ of the members with a precision of $80\%$.
(a) (a) CIFAR-10 (WRN28-10)
(b) (b) CIFAR-10 (VGG-16)
(c) (d) CIFAR-100 (DenseNet-121)
(d) (d) CIFAR-100 (SmallNet)
(e) (e) CINIC-10 (VGG-16)
(f) (f) Adult (MLP)
(g) (g) Credit (MLP)
@@ 4.2.3 Balanced accuracy and AUC
The balanced accuracy is the arithmetic mean of sensitivity and specificity, the higher the better. AUC value indicates the overall discriminatory power of the model over all possible TPR-FPR tradeoffs, the higher the better. Table IV shows the balanced Accuracy and AUC of all the MIA methods averaged out of 5 runs. The highest metric values and the metric values of LDC-MIA have been highlighted. LDC-MIA achieves the highest AUC values across all the datasets. Regarding balanced accuracy values, LDC-MIA has close if not better results compared to the best-performing MIAs.
@@ 4.2.4 Improvement on TPR at various hardness levels
4.3 Comparison with LiRA
LiRA [3] is a state-of-the-art MIA that can achieve high TPRs in the low FPR regions. To determine membership, LiRA calculates the likelihood ratio, which represents the ratio of the likelihood of a data point being in the target model’s training dataset to the likelihood of it being from a different, unknown dataset. The likelihood ratio is computed based on the output probabilities of $N$ shadow models and membership is determined based on the likelihood ratio. In our experiments, we compare LDC-MIA with online LiRA. WideResNet28-10 is the target model and is trained on the CIFAR-10 dataset. We then gradually increase the number of shadow or reference models used by LiRA and LDC-MIA and compare the TPRs at $1\%$ FPR as well as AUCs. The results are shown in Table V.
TABLE V: Performance of LiRA and LDC-MIA on CIFAR-10 (WRN28-10).
Auxiliary TPR at 1%FPR AUC
Models LiRA LDC-MIA LiRA LDC-MIA
N=2 2.38% 9.07% 0.599 0.794
N=4 3.21% 8.21% 0.676 0.806
N=8 5.46% 9.02% 0.656 0.802
N=16 7.12% 9.05% 0.688 0.807
N=32 15.81% 8.76% 0.757 0.805
N=64 18.33% 9.01% 0.771 0.804
N=128 20.75% 8.59% 0.781 0.801
N=256 21.32% 8.97% 0.792 0.806
From the results, we can see that the performance of LDC-MIA does not change significantly as the number of auxiliary models increases, while LiRA’s performance improves with more auxiliary models. However, the AUC of LDC-MIA outperforms the LiRA’s, with different numbers of auxiliary models, from 2 to 256. The better AUC across the board indicates that our method LDC-MIA, which uses multiple features, can help to distinguish members from non-members better on average. When the number of auxiliary models is no more than 16, the TPRs of LDC-MIA at an FPR of $1\%$ are higher than those of LiRA. Note that LiRA achieves better TPRs in the low FPRs when more than 32 auxiliary models are used; however, as discussed in the next section, this could lead to cost concerns.
4.4 Attack Cost
In real-world attacks, the cost of the attack is another key factor to consider alongside the attack performance. A cost-efficient approach, sometimes even a less performant one, may work better for some use cases. Let’s consider a few scenarios as follows. First, attackers have a limited budget for an MIA. Not having access to powerful computation hardware disallows them from training many shadow models or reference models for a high-cost attack. Second, MIA is used as an auditing tool to evaluate data privacy [3]. where a model provider needs to evaluate many of the training data to ensure a sufficiently low likelihood of privacy leaks before releasing the model. For this use case, an MIA approach whose cost is sub-linear to the number of sample data is preferred. Third, the aforementioned model provider might need to frequently update the model training dataset for retraining [11, 24]. Running MIA for data leakage audit every time after retraining asks for a balanced tradeoff between the utility and the cost of the MIA approach. Lastly, a less costly MIA approach may serve better with a tight timeline. For example, a model provider must release an updated model by a deadline and finish the MIA data leakage audit on time. A less costly MIA usually requires less computing and thus is faster.
We compare the cost of MIAs by comparing the cost of computing and the cost of data. We approximate the compute cost by calculating the training time and inference time of all auxiliary models used in the MIA (including shadow and reference models). Note that all computing experiments run on the same hardware configuration. We approximate the data cost of different MIAs as the number of auxiliary models used because the MIAs we evaluate take a similar amount of training data for auxiliary models. Therefore, the more auxiliary models are used, the larger the data cost for the MIA. Both LDC-MIA and Salem et al. [29] need to train a neural network-based MIA classifier in addition to the auxiliary models. These two classifiers employ a shallow multilayer perception (MLP) architecture. It takes around 22 seconds to train such a classifier in all the datasets we experimented with. This is a significantly small cost compared to training one shadow model for image classification. Therefore, these MIA classifiers’ training and inference time are not included in the discussion below, where we use the task of attacking CIFAR-10 as an example to demonstrate the attack costs of different MIAS. Based on the results in Section 4.3, LDC-MIA does not have much performance gains from employing additional auxiliary models. Hence, we calculate the cost of LDC-MIA by including only one shadow model and one reference model.
Table VI shows the costs of attacking WideResNet28-10 on the CIFAR-10 dataset.
TABLE VI: The attack costs of different MIAs.
Attack Auxiliary TPR at AUC Training cost Inference cost (s)
Method Models (N) $1\%$FPR (minutes) per attack
Salem et al. [29] 1 2.5% 0.731 19 24
Yoem et al. [39] - 0.0% 0.723 - 19
Watson et al. [37] 1 3.3% 0.707 19 38
Watson et al. [37] 10 3.7% 0.742 189 221
LiRA[3] 16 7.12% 0.688 304 348
LiRA[3] 256 21.32% 0.792 4867 4942
Ours 2 9.1% 0.794 39 43
We perform the MIAs on 25k target samples and show the sum of attack costs of all the samples. The training cost of the attack does not change as the number of target samples to attack increases, except for LiRA. The training cost for LiRA increases linearly as the number of target samples changes because LiRA requires training multiple shadow models for each target sample for a reliable estimate of the likelihood ratio.
Yeom et al. [39] has the least cost among all, as it requires no auxiliary models and only needs one-time inference on the target model per target sample. Coming with the low cost are the low AUC and TPR of this approach. Our method LDC-MIA has a similar cost to Salem et al. [29]. Both attacks require training a few auxiliary models that, once ready, can be used for all target sample attacks. Then, per target sample, the attack takes one or two inferences on the target model. The total cost is in the order of tens of minutes. Watson et al. [37] method can opt-in to use multiple reference models for better attack performance, though with higher training cost, which is linear to the number of the auxiliary models, as illustrated in Table VI.
In contrast, for LiRA, the training cost is linear to the product of the number of auxiliary models and the number of target samples. For every target sample, LiRA must train a different set of auxiliary models tailored to that particular sample. This is different from all other MIAs, including LDC-MIA. Multiple target samples can reuse the same auxiliary models in all other MIAs. Thus, the training cost can be amortized across a batch of target sample attacks.